# Shape-constrained density estimation with Wasserstein projection

### 09 Mar 2026 v1

@@ 3. Monotone density estimation
Let $p=2$ and let $\mu=\delta_{1}$ be the point mass at $1$. For $b>0$, let $\nu_{b}=\mathrm{Unif}(0,b)$ be the uniform distribution on $[0,b]$. A straightforward computation shows that
@@
We use the first-order condition in Corollary 2.9. By Proposition 3.4, $\operatorname{proj}_{\mathcal{F}_{m,2}}\mu$ is compactly supported. It suffices to test with $\nu\in\mathcal{F}_{2,m}$ which is compactly supported. Then $Q=Q_{\nu}:[0,1]\rightarrow\mathbb{R}_{+}$ is convex, non-decreasing and $Q(0)=\lim_{u\downarrow 0}Q(u)=0$. In particular, $Q$ is bounded. Note that $Q_{\mu}(u)=\frac{3}{2}u$. We claim that
(3.4) $\begin{split}&\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}\right|_{\epsilon=0^{+}}\int_{0}^{1}\left(\frac{3}{2}u+\epsilon\left(Q(u)-\frac{3}{2}u\right)-1\right)^{2}\mathrm{d}u\\ &=\int_{0}^{1}\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}\right|_{\epsilon=0^{+}}\left(\frac{3}{2}u+\epsilon\left(Q(u)-\frac{3}{2}u\right)-1\right)^{2}\mathrm{d}u\\ &=\int_{0}^{1}2\left(\frac{3}{2}u-1\right)\left(Q(u)-\frac{3}{2}u\right)\mathrm{d}u=\int_{0}^{1}(3u-2)Q(u)\mathrm{d}u\geq 0.\end{split}$
@@
$\begin{split}\int_{0}^{1}(3u-1)Q(u)\mathrm{d}u&=\int_{0}^{1}(3u-2)\left(\int_{0}^{u}q(s)\mathrm{d}s\right)\mathrm{d}u\\ &=\int_{0}^{1}q(s)\left(\int_{s}^{1}(3u-2)\mathrm{d}u\right)\mathrm{d}s\\ &=\int_{0}^{1}q(s)K(s)\mathrm{d}s,\end{split}$
where $K(s)=\frac{1}{2}(3s-1)(1-s)$, $s\in[0,1]$. Note that $K(s)\leq 0$ on $0,\frac{1}{3}]$ and is positive on $(\frac{1}{3},1]$. Moreover, one may verify that $\int_{0}^{1}K(s)\mathrm{d}s=0$. Since $q\geq 0$ is non-decreasing, we have
$\begin{split}\int_{0}^{1}(3u-1)Q(u)\mathrm{d}u&=\int_{0}^{\frac{1}{3}}q(s)K(s)\mathrm{d}s+\int_{\frac{1}{3}}^{1}q(s)K(s)\mathrm{d}s\\ &\geq\int_{0}^{\frac{1}{3}}q\Big(\frac{1}{3}\Big)K(s)\mathrm{d}s+\int_{\frac{1}{3}}^{1}q\Big(\frac{1}{3}\Big)K(s)\mathrm{d}s\\ &=q\Big(\frac{1}{3}\Big)\int_{0}^{1}K(s)\mathrm{d}s=0.\end{split}$
This proves the claim. By Corollary 2.9, we conclude that $\operatorname{proj}_{\mathcal{F}_{m,2}}\mu=\mathrm{Unif}(0,\frac{3}{2})$.
Here is the main result of this section.
@@
Let $p=2$ and $\mu_{n}=\sum_{i=1}^{n}\frac{1}{n}\delta_{x_{i}}$ where $x_{1},\ldots,x_{n}>0$. Then, the density of $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{m,2}}\mu_{n}$, which is non-increasing, is compactly supported and piecewise constant with finitely many pieces.
@@
We will argue in terms of the quantile functions. Let $Q_{0}:=Q_{\mu_{n}}$ be the quantile function of the data. Note that $Q_{0}>0$ and is piecewise constant. By Proposition 3.4, $\hat{\mu}_{n}$ has a non-increasing density which is compactly supported on $\mathbb{R}_{+}$. Let $\hat{Q}:=Q_{\hat{\mu}_{n}}$ be its quantile function. To prove the theorem, we need to show that $\hat{Q}$ is piecewise affine.
@@
(3.6) $\int_{0}^{1}h(s)\phi(s)\mathrm{d}s\leq\int_{0}^{1}\hat{h}(s)\phi(s)\mathrm{d}s,\quad h\in\mathcal{H}.$
@@
Now since $\mathcal{H}$ is a cone, the supremum must be 0. Thus, we have
@@
Next, consider the identity
$\int_{0}^{1}\hat{h}(s)\phi(s)\mathrm{d}s=0.$
Integrating by parts, we have
@@
We claim that $Z$ is finite. If so, $\hat{h}$ can only increase (jump) at finitely many points. Thus, $\hat{Q}(u)=\int_{0}^{u}\hat{h}(s)\mathrm{d}s$ is piecewise affine, implying that the density of $\hat{\mu}$ is piecewise constant (with finitely many pieces). We prove this claim in two steps.
@@
Since $\hat{Q}$ is convex and bounded, it is continuous. Also, $Q_{0}(u)$ is piecewise constant. Thus, the left and right second derivatives $\Phi^{\prime\prime}(t-)$ and $\Phi^{\prime\prime}(t+)$ exist everywhere on $(0,1)$. Moreover, they agree (so $\Phi^{\prime\prime}(t)$ exists) and are equal to $Q_{0}(t)-\hat{Q}(t)$, except at finitely many points where $Q_{0}$ jumps. (There can be none if $Q_{0}$ is constant.)
@@ 4. Log-concave density estimation
    $Q$ is absolutely continuous and there exists a version of $Q^{\prime}$ such that $Q^{\prime}>0$ and $\frac{1}{Q^{\prime}}$ is concave on $(0,1)$.
@@
for some interval $(a,b)$ and density $f$ such that $f>0$ on $(a,b)$ and $\phi:=-\log f$ is convex. In particular, $f$ is continuous on $(a,b)$ (see [28, Theorem 10.1]). By [6, Proposition A.18], for any $0<u_{0}<u_{1}<1$, we have
@@
Since $\phi$ is convex, its one-sided derivatives $\phi^{\prime}(x-)$ and $\phi^{\prime}(x+)$ exist for all $x\in(a,b)$. It follows that $\frac{1}{Q^{\prime}}$ is one-sided differentiable, and we have
@@
Since $\phi^{\prime}(Q(u)+)\leq\phi^{\prime}(Q(v)-)$ for $u<v$ from the convexity of $\phi$, we have that $\frac{1}{Q^{\prime}}$ is concave.
@@
Since $h^{\prime}$ is monotone, $\log f$ is concave and the theorem is proved. ∎
Next, we show that $\mathcal{F}_{lc}$ is $\mathcal{W}_{p}$-closed and displacement convex. Hence, the Wasserstein projection $\operatorname{proj}_{\mathcal{F}_{lc}}^{(p)}$ is well-defined. We stress that displacement convexity of $\mathcal{F}_{lc}$ only holds in the univariate setting. It was proved in [33] that the space of log-concave distributions on $\mathbb{R}^{d}$ is not displacement convex when $d\geq 2$.
@@
The set $\mathcal{F}_{lc}\subset\mathcal{P}(\mathbb{R})$ of log-concave distributions on $\mathbb{R}$ is displacement convex [^5] From [33, Footnote 1], this result was proved by Young-Heon Kim but his proof is not publicly available. For completeness we provide a self-contained proof using Lemma 4.1. and is closed in $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ for any $p\in[1,\infty)$.
@@
We first observe that $\mathcal{F}_{lc}$ is closed in $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ for $p\in[1,\infty)$. Suppose $\mu_{n}\in\mathcal{F}_{lc}$ and $\mathcal{W}_{p}(\mu_{n},\mu)\rightarrow 0$ for some $\mu\in\mathcal{P}_{p}(\mathbb{R})$. Then, $\mu_{n}$ converges weakly to $\mu$ by [39, Theorem 6.9]. Since $\mathcal{F}_{lc}$ is closed under weak convergence (see [34, Proposition 3.6]), we have $\mu\in\mathcal{F}_{lc}$.
@@
Let $\mu=\frac{1}{2}\delta_{-1}+\frac{1}{2}\delta_{1}$. We claim that $\operatorname{proj}_{\mathcal{F}_{lc}}\mu=\mathrm{Unif}(-3/2,3/2)$.
Since $\mu$ is symmetric about 0 and $\mathcal{F}_{lc}$ is invariant under inversion (that is, $T_{\#}\mathcal{F}_{lc}=\mathcal{F}_{lc}$ where T(x)=xT(x)=-x), by Proposition 2.11 we have that $\operatorname{proj}_{\mathcal{F}_{lc}}\mu$ is symmetric about 0. Conditioning on $\mathbb{R}_{+}$, we see that the projection problem $\operatorname{proj}_{\mathcal{F}_{lc}}\mu$ is equivalent to $\operatorname{proj}_{\{\delta_{0}\}\cup\tilde{\mathcal{F}}}\delta_{1}$, where $\tilde{\mathcal{F}}$ is the collection of log-concave and non-increasing densities on $\mathbb{R}_{+}$. It is easily seen (see Proposition 4.6(i) below) that $\delta_{0}$ is not optimal. From Example 3.5, $\mathrm{Unif}(0,\frac{3}{2})$ minimizes $\nu\mapsto\mathcal{W}_{2}(\nu,\delta_{1})$ over all non-increasing densities on $\mathbb{R}_{+}$, log concave or not. Since $\mathrm{Unif}(0,\frac{3}{2})$ is log concave, it must be optimal for $\operatorname{proj}_{\tilde{\mathcal{F}}}\delta_{1}$. It follows that $\operatorname{proj}_{\mathcal{F}_{lc}}\mu=\mathrm{Unif}(-\frac{3}{2},\frac{3}{2})$.
@@
Let the true distribution $\mu^{*}$ be an element of $\mathcal{F}_{lc}$. Consider the Wasserstein projection estimator $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{lc}}\mu_{n}$ where $p=2$ and $\mu_{n}=\sum_{i=1}^{n}\frac{1}{n}\delta_{X_{i}}$ is the empirical measure of $n$ independent samples from $\mu^{*}$.
@@
Next, we prove analogues of Proposition 3.4 and Theorem 3.6 in the log-concave setting.
@@
(4.6) $\begin{split}\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}\mathcal{W}_{2}^{2}(\nu_{\epsilon},\mu)\right|_{\epsilon=0^{+}}&=\left.\int_{0}^{1}2(-\epsilon+2\epsilon u-Q(u))(2u-1)\mathrm{d}u\right|_{\epsilon=0^{+}}\\ &=-4\int_{0}^{1}uQ(u)\mathrm{d}u,\end{split}$
@@
(ii) Let $\mu$ be supported on a compact interval $[a,b]$. Suppose the support of $\hat{\mu}$ is unbounded on the right. That is, the density $\hat{f}$ of $\hat{\mu}$ (which exists since $\hat{\mu}$ is not a point mass) is positive on $x,\infty)$ for all $x$ sufficiently large. We will construct $\tilde{\mu}\in\mathcal{F}_{lc}$ such that its support is bounded on the right and $\mathcal{W}_{2}(\tilde{\mu},\mu)<\mathcal{W}_{2}(\hat{\mu},\mu)$. The left tail, if unbounded, can be handled similarly.
Let $T:\mathbb{R}\rightarrow[a,b]$ be the Brenier map, which is $\hat{\mu}-$a.e. unique, that pushforwards $\hat{\mu}$ to $\mu$: $T_{\#}\hat{\mu}=\mu$. For $R>b$ sufficiently large such that $Z_{R}:=\hat{\mu}((-\infty,R])>0$, consider the probability measure $\tilde{\mu}_{R}$, which is concentrated on $(-\infty,R]$, defined by
$\mathrm{d}\tilde{\mu}_{R}(x):=\frac{1}{Z_{R}}I_{(-\infty,R]}(x)\mathrm{d}\hat{\mu}(x).$
Note that $\tilde{\mu}_{R}$ is log-concave since its density is
$\tilde{f}_{R}(x)=\frac{1}{Z_{R}}I_{(-\infty,R]}\hat{f}(x).$
Let $T_{R}:=\min\{T,R\}$ and note that $(T_{R})_{\#}\tilde{\mu}_{R}=\mu$. Hence, we have
$\begin{split}\mathcal{W}_{2}^{2}(\tilde{\mu}_{R},\mu)&\leq\int_{-\infty}^{R}(x-T_{R}(x))^{2}\mathrm{d}\tilde{\mu}_{R}(x)\\ &=\frac{1}{Z_{R}}\int_{-\infty}^{R}(x-T(x))^{2}\mathrm{d}\hat{\mu}(x)=:\frac{1}{Z_{R}}A_{R}.\end{split}$
On the other hand, we have
$\mathcal{W}_{2}^{2}(\hat{\mu},\mu)=\int_{-\infty}^{R}(x-T(x))^{2}\mathrm{d}\hat{\mu}(x)+\int_{R}^{\infty}(x-T(x))^{2}\mathrm{d}\hat{\mu}(x)=:A_{R}+B_{R}.$
Consider the difference
$\mathcal{W}_{2}^{2}(\tilde{\mu}_{R},\mu)-\mathcal{W}_{2}^{2}(\hat{\mu},\mu)=\frac{1-Z_{R}}{Z_{R}}A_{R}-B_{R}.$
Since $Z_{R}\uparrow 1$ and $A_{R}\uparrow\mathcal{W}_{2}^{2}(\hat{\mu},\mu)$, for $R$ sufficiently large we have
$\frac{1-Z_{R}}{Z_{R}}A_{R}\leq 2(1-Z_{R})\mathcal{W}_{2}^{2}(\hat{\mu},\mu).$
On the other hand, since $T(x)\leq b$, for $R>b$ we have
$B_{R}=\int_{R}^{\infty}(x-T(x))^{2}\mathrm{d}\hat{\mu}(x)\geq(R-b)^{2}\int_{R}^{\infty}\mathrm{d}\hat{\mu}(x)=(R-b)^{2}(1-Z_{R}).$
It follows that
$\mathcal{W}_{2}^{2}(\tilde{\mu}_{R},\mu)-\mathcal{W}_{2}^{2}(\hat{\mu},\mu)\leq(1-Z_{R})(2\mathcal{W}_{2}^{2}(\hat{\mu},\mu)-(R-b)^{2}),$
which is negative when $R$ is sufficiently large. [^6] The log-concave density $\hat{f}$ decays at least exponentially. This implies that $1-Z_{R}=O(e^{-\alpha R})$ for some constant $\alpha>0$. We can take $\tilde{\mu}=\tilde{\mu}_{R}$ for such an $R$. ∎
@@
Let $p=2$ and $\mu_{n}=\sum_{i=1}^{n}\frac{1}{n}\delta_{x_{i}}$ where $x_{1},\ldots,x_{n}\in\mathbb{R}$, be an empirical measure, and let $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{lc}}\mu_{n}$. If $\mu_{n}$ is a point mass, then $\hat{\mu}_{n}=\mu_{n}$. Otherwise, $\hat{\mu}_{n}$ has a compactly supported density which is piecewise log-affine with finitely many pieces.
@@
where $F$ is the distribution function of $\mu$. Since $F$ is continuous and strictly increasing on the range of $Q_{\mu}$, we have
[…]
@@
Since $\mu_{n}$ is not a point mass, by Proposition 4.6, $\hat{\mu}_{n}$ has a compactly supported density which is log-concave. Let $\hat{Q}:=Q_{\hat{\mu}_{n}}$ be the quantile function of $\hat{\mu}_{n}$, and let $\hat{h}:=\frac{1}{\hat{Q}^{\prime}}$ which is positive and concave. Since $\hat{Q}$ is bounded, we have that $\frac{1}{\hat{h}}\in L^{1}(0,1)$. From Lemma 4.8, the theorem is proved if we show that $\hat{h}$ is piecewise affine (with finitely many pieces).
@@
$Q(u)=a+\int_{0}^{u}\frac{1}{h(s)}\mathrm{d}s$
where $a\in\mathbb{R}$ and $h>0$ is concave and $\frac{1}{h}\in L^{1}(0,1)$ (we let $\mathcal{H}$ denote the set of such functions). We obtain
@@
whenever the derivative can be taken inside the integral sign.
@@
We consider the former case (the other one is similar).

### 10 Apr 2026 v2

@@ 3. Monotone density estimation
Let $p=2$ and let $\mu=\delta_{1}$ be the point mass at $1$. For $b>0$, let $\nu_{b}=\mathrm{Unif}(0,b)$ be the uniform distribution on $[0,b]$. Then
@@
We use the first-order condition in Corollary 2.9. By Proposition 3.4, $\operatorname{proj}_{\mathcal{F}_{m,2}}\mu$ is compactly supported. It suffices to test with $\nu\in\mathcal{F}_{2,m}$ which is compactly supported. Then $Q=Q_{\nu}:[0,1]\rightarrow\mathbb{R}_{+}$ is convex, non-decreasing and $Q(0)=\lim_{u\downarrow 0}Q(u)=0$. In particular, $Q$ is bounded. Note that $Q_{\mu}(u)=(3/2)u$. We claim that
(3.5) $\begin{split}&\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}\right|_{\epsilon=0^{+}}\int_{0}^{1}\left(\frac{3}{2}u+\epsilon\left(Q(u)-\frac{3}{2}u\right)-1\right)^{2}\mathrm{d}u\\ &=\int_{0}^{1}2\left(\frac{3}{2}u-1\right)\left(Q(u)-\frac{3}{2}u\right)\mathrm{d}u=\int_{0}^{1}(3u-2)Q(u)\mathrm{d}u\geq 0.\end{split}$
@@
$\begin{split}\int_{0}^{1}(3u-2)Q(u)\mathrm{d}u&=\int_{0}^{1}(3u-2)\left(\int_{0}^{u}q(s)\mathrm{d}s\right)\mathrm{d}u\\ &=\int_{0}^{1}q(s)\left(\int_{s}^{1}(3u-2)\mathrm{d}u\right)\mathrm{d}s\\ &=\int_{0}^{1}q(s)K(s)\mathrm{d}s,\end{split}$
where K(s)=(3s1)(1s)/2K(s)=(3s-1)(1-s)/2. Note that $K(s)$ is non-positive on $[0,1/3]$, positive on $(1/3,1]$, and integrates to 0 on $[0,1]$. Since $q\geq 0$ is non-decreasing, we have
$\begin{split}\int_{0}^{1}(3u-1)Q(u)\mathrm{d}u&=\int_{0}^{1/3}q(s)K(s)\mathrm{d}s+\int_{1/3}^{1}q(s)K(s)\mathrm{d}s\\ &\geq\int_{0}^{1/3}q\Big(\frac{1}{3}\Big)K(s)\mathrm{d}s+\int_{1/3}^{1}q\Big(\frac{1}{3}\Big)K(s)\mathrm{d}s\\ &=q\Big(\frac{1}{3}\Big)\int_{0}^{1}K(s)\mathrm{d}s=0.\end{split}$
This proves the claim, and we conclude that $\operatorname{proj}_{\mathcal{F}_{m,2}}\mu=\mathrm{Unif}(0,3/2)$.
We arrive at the main result of this section.
@@
Let $p=2$ and $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{x_{i}}$ where $x_{1},\ldots,x_{n}>0$. Then, the density of $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{m,2}}\mu_{n}$, which is non-increasing, is compactly supported and piecewise constant with finitely many pieces.
@@
We will argue in terms of the quantile functions. Let $Q_{0}:=Q_{\mu_{n}}$ be the quantile function of the data. Note that $Q_{0}>0$ and is piecewise constant. By Proposition 3.4, $\hat{\mu}_{n}$ has a non-increasing density which is compactly supported on $\mathbb{R}_{+}$. Let $\hat{Q}:=Q_{\hat{\mu}_{n}}$ be its quantile function. To prove the theorem, we need to show that $\hat{Q}$ is piecewise affine (see (3.2)).
@@
$\int_{0}^{1}h(s)\phi(s)\mathrm{d}s\leq\int_{0}^{1}\hat{h}(s)\phi(s)\mathrm{d}s,\quad h\in\mathcal{H}.$
@@
Since $\mathcal{H}$ is a cone, the supremum must be 0. Thus, we have
@@
Next, consider the identity $\int_{0}^{1}\hat{h}(s)\phi(s)\mathrm{d}s=0$. Integrating by parts, we have
@@
We claim that $Z$ is finite. If so, $\hat{h}$ can only increase (jump) at finitely many points. Thus, $\hat{Q}(u)=\int_{0}^{u}\hat{h}(s)\mathrm{d}s$ is piecewise affine, implying that the density of $\hat{\mu}$ is piecewise constant with finitely many pieces. We prove this claim in two steps.
@@
Since $\hat{Q}$ is convex and bounded, it is continuous. Also, $Q_{0}(u)$ is piecewise constant. Thus, the left and right second derivatives $\Phi^{\prime\prime}(t-)$ and $\Phi^{\prime\prime}(t+)$ exist everywhere on $(0,1)$. Moreover, they agree (so $\Phi^{\prime\prime}(t)$ exists) and are equal to $Q_{0}(t)-\hat{Q}(t)$, except at finitely many points where $Q_{0}$ jumps. [^7] There can be none if $Q_{0}$ is constant. In this case, $\hat{\mu}_{n}$ is a uniform distribution by Example 3.5 (and equivariance under scaling).
@@ 4. Log-concave density estimation
    $Q$ is absolutely continuous and there exists a version of $Q^{\prime}$ such that $Q^{\prime}>0$ and $1/Q^{\prime}$ is concave on $(0,1)$.
@@
for some interval $(a,b)$ and density $f$ such that $f>0$ on $(a,b)$ and $\phi:=-\log f$ is convex. In particular, $f$ is continuous on $(a,b)$ (see [29, Theorem 10.1]). By [6, Proposition A.18], for any $0<u_{0}<u_{1}<1$, we have
@@
Since $\phi$ is convex, its one-sided derivatives $\phi^{\prime}(x-)$ and $\phi^{\prime}(x+)$ exist for all $x\in(a,b)$. It follows that $1/Q^{\prime}$ is one-sided differentiable, and we have
@@
Since $\phi^{\prime}(Q(u)+)\leq\phi^{\prime}(Q(v)-)$ for $u<v$ from the convexity of $\phi$, we have that $1/Q^{\prime}$ is concave.
@@
Since $h^{\prime}$ is non-increasing, $\log f$ is concave and the theorem is proved. ∎
Next, we show that $\mathcal{F}_{lc}$ is $\mathcal{W}_{p}$-closed and displacement convex for any $p>1$. Hence, the Wasserstein projection $\operatorname{proj}_{\mathcal{F}_{lc}}^{(p)}$ is well-defined. We stress that displacement convexity of $\mathcal{F}_{lc}$ only holds in the univariate setting. It was proved in [34] that the space of log-concave distributions on $\mathbb{R}^{d}$ is not displacement convex when $d\geq 2$.
@@
The set $\mathcal{F}_{lc}\subset\mathcal{P}(\mathbb{R})$ of log-concave distributions on $\mathbb{R}$ is displacement convex [^8] From [34, Footnote 1], this result was proved by Young-Heon Kim but his proof is not publicly available. For completeness we provide a self-contained proof using Lemma 4.1. and is closed in $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ for any $p\in[1,\infty)$.
@@
We first observe that $\mathcal{F}_{lc}$ is closed in $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ for $p\in[1,\infty)$. Suppose $\mu_{n}\in\mathcal{F}_{lc}$ and $\mathcal{W}_{p}(\mu_{n},\mu)\rightarrow 0$ for some $\mu\in\mathcal{P}_{p}(\mathbb{R})$. Then, $\mu_{n}$ converges weakly to $\mu$ by [40, Theorem 6.9]. Since $\mathcal{F}_{lc}$ is closed under weak convergence (see [35, Proposition 3.6]), we have $\mu\in\mathcal{F}_{lc}$.
@@
Let $\mu=(1/2)\delta_{-1}+(1/2)\delta_{1}$. We claim that $\operatorname{proj}_{\mathcal{F}_{lc}}\mu=\mathrm{Unif}(-3/2,3/2)$.
Since $\mu$ is symmetric about 0 and $\mathcal{F}_{lc}$ is invariant under inversion (that is, $T_{\#}\mathcal{F}_{lc}=\mathcal{F}_{lc}$ where T(x)=xT(x)=-x), by Proposition 2.11 we have that $\operatorname{proj}_{\mathcal{F}_{lc}}\mu$ is symmetric about 0. Conditioning on $\mathbb{R}_{+}$, we see that the projection problem $\operatorname{proj}_{\mathcal{F}_{lc}}\mu$ is equivalent to $\operatorname{proj}_{\{\delta_{0}\}\cup\tilde{\mathcal{F}}}\delta_{1}$, where $\tilde{\mathcal{F}}$ is the collection of log-concave and non-increasing densities on $\mathbb{R}_{+}$. It is easily seen (see Proposition 4.6(i) below) that $\delta_{0}$ is not optimal. From Example 3.5, $\mathrm{Unif}(0,3/2)$ minimizes $\nu\mapsto\mathcal{W}_{2}(\nu,\delta_{1})$ over all non-increasing densities on $\mathbb{R}_{+}$, log concave or not. Since $\mathrm{Unif}(0,3/2)$ is log concave, it must be optimal for $\operatorname{proj}_{\tilde{\mathcal{F}}}\delta_{1}$. It follows that $\operatorname{proj}_{\mathcal{F}_{lc}}\mu=\mathrm{Unif}(-3/2,3/2)$.
@@
Let the true distribution $\mu^{*}$ be an element of $\mathcal{F}_{lc}$. Consider the Wasserstein projection estimator $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{lc}}\mu_{n}$ where $p=2$ and $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{X_{i}}$ is the empirical measure of $n$ independent samples from $\mu^{*}$.
@@
Next, we prove the analogues of Proposition 3.4 and Theorem 3.6 in the log-concave setting.
@@
(4.6) $\begin{split}\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}\mathcal{W}_{2}^{2}(\nu_{\epsilon},\mu)\right|_{\epsilon=0^{+}}&=-4\int_{0}^{1}uQ(u)\mathrm{d}u,\end{split}$
@@
(ii) We prove the following one-sided statement which is slightly stronger. Suppose that $\mu$ is supported on $(-\infty,b]$ for some $b<\infty$, so that $Q_{\mu}\leq b$. We show that the support of $\hat{\mu}=\operatorname{proj}_{\mathcal{F}_{lc}}\mu$ is also bounded on the right. That is, if $\hat{Q}:=Q_{\hat{\mu}}$ then $\sup_{u\in(0,1)}\hat{Q}(u)<\infty$. The left tail can be handled similarly.
Suppose on the contrary that $\sup_{u\in(0,1)}\hat{Q}(u)=\infty$, so the right tail of the density of $\hat{\mu}$ is unbounded on the right. We will construct $\tilde{\mu}\in\mathcal{F}_{lc}$, with a bounded right tail, such that $\mathcal{W}_{2}(\tilde{\mu},\mu)\leq\mathcal{W}_{2}(\hat{\mu},\mu)$. Since $\sup_{u\in(0,1)}\hat{Q}(u)=\infty$, for $R>b$ sufficiently large we have $Z_{R}:=\hat{\mu}((-\infty,R])>0$. Note that $Z_{R}\uparrow 1$ as $R\uparrow\infty$. When $Z_{R}>0$, let $\tilde{\mu}_{R}$ be $\hat{\mu}$ conditioned on $(-\infty,Z_{R}]$, and note that $\tilde{\mu}_{R}$ is log-concave. Its quantile function is given by
$Q_{\tilde{\mu}_{R}}(u)=\hat{Q}(Z_{R}u),\quad u\in(0,1).$
Note that $\tilde{Q}_{R}\leq\hat{Q}(Z_{R})<\infty$.
We estimate $\mathcal{W}_{2}^{2}(\hat{\mu},\mu)-\mathcal{W}_{2}^{2}(\tilde{\mu}_{R},\mu)$ from below. Write
$\begin{split}&\mathcal{W}_{2}^{2}(\hat{\mu},\mu)-\mathcal{W}_{2}^{2}(\tilde{\mu}_{R},\mu)\\ &=\int_{0}^{1}(\hat{Q}(u)-Q_{\mu}(u))^{2}\mathrm{d}u-\int_{0}^{1}(\hat{Q}(Z_{R}u)-Q_{\mu}(u))^{2}\mathrm{d}u\\ &=\int_{0}^{1}(\hat{Q}(u)-\hat{Q}(Z_{R}u))(\hat{Q}(u)+\hat{Q}(Z_{R}u)-2Q_{\mu}(u))\mathrm{d}u\\ &\geq\int_{0}^{1}(\hat{Q}(u)-\hat{Q}(Z_{R}u))(\hat{Q}(u)+\hat{Q}(Z_{R}u)-2b)\mathrm{d}u,\end{split}$
where in the last line we used the inequalities $\hat{Q}(u)-\hat{Q}(Z_{R}u)\geq 0$ ($\hat{Q}$ is non-decreasing and $u\geq Z_{R}u$) and $Q_{\mu}\leq b$.
Expanding and rearranging the integrals, we see that the above is equal to
$\begin{split}&\int_{0}^{1}\hat{Q}^{2}(u)\mathrm{d}u-\int_{0}^{1}\hat{Q}^{2}(Z_{R}u)\mathrm{d}u-2b\left(\int_{0}^{1}\hat{Q}(u)\mathrm{d}u-\int_{0}^{1}\hat{Q}(Z_{R}u)\mathrm{d}u\right)\\ &=\int_{0}^{1}\hat{Q}^{2}(u)\mathrm{d}u-\frac{1}{Z_{R}}\int_{0}^{Z_{R}}\hat{Q}^{2}(u)\mathrm{d}u-2b\left(\int_{0}^{1}\hat{Q}(u)\mathrm{d}u-\frac{1}{Z_{R}}\int_{0}^{Z_{R}}\hat{Q}(u)\mathrm{d}u\right)\\ &=\int_{Z_{R}}^{1}\left(\hat{Q}^{2}(u)-2b\hat{Q}(u)\right)\mathrm{d}u+\frac{Z_{R}-1}{Z_{R}}\int_{0}^{Z_{R}}\left(\hat{Q}^{2}(u)-2b\hat{Q}(u)\right)\mathrm{d}u.\end{split}$
Since $\hat{Q}(u)\geq R$ for $u\geq Z_{R}$, when $R\geq\max\{0,2b\}$ the first term is bounded below by $R(R-2b)(1-Z_{R})$. On the other hand, for $R$ sufficiently large such that $Z_{R}\leq\frac{1}{2}$, the second term is bounded above by
$2(1-Z_{R})K,\quad K:=\int_{0}^{1}\left(\hat{Q}^{2}(u)-2|b||\hat{Q}(u)|\right)\mathrm{d}u<\infty.$
(Recall that $\hat{Q}\in L^{2}([0,1])$. It follows that for $R$ sufficiently large, we have
$\begin{split}\mathcal{W}_{2}^{2}(\hat{\mu},\mu)-\mathcal{W}_{2}^{2}(\tilde{\mu}_{R},\mu)&\geq R(R-b)(1-Z_{R})-2(1-Z_{R})K\\ &=(1-Z_{R})(R(R-b)-2K),\end{split}$
which is positive when additionally R(Rb)2K>0R(R-b)-2K>0. Thus we may let $\tilde{\mu}=\tilde{\mu}_{R}$ for such a value of $R$. ∎
@@
Let $p=2$ and $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{x_{i}}$ where $x_{1},\ldots,x_{n}\in\mathbb{R}$, be an empirical measure, and let $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{lc}}\mu_{n}$. If $\mu_{n}$ is a point mass, then $\hat{\mu}_{n}=\mu_{n}$. Otherwise, $\hat{\mu}_{n}$ has a compactly supported density which is piecewise log-affine with finitely many pieces.
@@
where $F$ is the distribution function of $\mu$. Since $F$ is continuous and strictly increasing on the range of $Q_{\mu}$, the proof is completed by observing that the following statements are equivalent:
  • (i)
    $f$ is piecewise log-affine.
  • (ii)
    $\log f$ is piecewise affine.
  • (iii)
    $(\log f)^{\prime}$ is piecewise constant.
  • (iv)
    $h^{\prime}$ is piecewise constant
  • (v)
    $h$ is piecewise affine.
@@
Since $\mu_{n}$ is not a point mass, by Proposition 4.6, $\hat{\mu}_{n}$ has a compactly supported density which is log-concave. Let $\hat{Q}:=Q_{\hat{\mu}_{n}}$ be the quantile function of $\hat{\mu}_{n}$, and let $\hat{h}:=1/\hat{Q}^{\prime}$ which is positive and concave. Since $\hat{Q}$ is bounded, we have that $1/h\in L^{1}(0,1)$. From Lemma 4.8, the theorem is proved if we show that $\hat{h}$ is piecewise affine with finitely many pieces.
@@
$Q(u)=a+\int_{0}^{u}\frac{1}{h(s)}\mathrm{d}s,$
where $a\in\mathbb{R}$ and $h>0$ is concave and $1/h\in L^{1}([0,1])$. Let $\mathcal{H}$ denote the set of such functions. We obtain
@@
whenever the derivative can be taken under the integral sign.
@@
We consider the former case; the other one is similar.